OpenStudy (anonymous):

With the functions: f(x) = x^2 + 3x and g(x) = -3 + x what is f(g(x)) ? Can anyone help with this? I keep getting x^2 + 3x - 9 but that isn't one of the options. The choices are: a. x^2 - 3x b. x^2 - 3x + 9 c. x^2 + 3x + 9 d. 3x + 9

OpenStudy (amistre64):

f is currently composed of x, we want to compose it if g(x), so let x=g(x) f(g(x)) = [g(x)]^2+3[g(x)] but we know g(x) is just a name for: x-3 soo f(g(x)) = (x-3)^2+3(x-3) and simplify as needed

OpenStudy (anonymous):

Aah, thank you! @amistre64 So the correct answer is just x^2 - 3x?

OpenStudy (amistre64):

hmm lets test it out (x-3)(x-3+3) (x-3)(x) i like it

OpenStudy (anonymous):

Thank you! <3